There are 50,000 people at a stadium watching a soccer match, and 40,000 of them are male. If 3 people are chosen at random, what is the probability that all 3 of them are male?
Answer by stanbon(75887)

There are 50,000 people at a stadium watching a soccer match, and 40,000 of them are male. If 3 people are chosen at random, what is the probability that all 3 of them are male?
-----
Binomial Problem with n = 3 and p = 4/5
---
P(x = 3) = (4/5)^3 = 64/125 = 0.512
=====================
Cheers,
Stan H.
=============